Separate this redox reaction into its balanced component half-reactions.

Cl2+2Na---->2NaCl

Oxidation of half reaction:


Recuction half reaction:
Answer #1
Concepts and reason

Cell potential is the potential difference between the two half cells in an electrochemical cell. The potential difference is created by the movement of electrons from one half cell to another.

The given electrochemical cell contains two half cells which adds up to give the overall balanced reaction.

Fundamentals

An electrochemical cell consists of two half cells. In one half, oxidation (loss of electron) occurs and in other half, reduction (gain of electron) occurs.

The oxidation half-cell is termed as anode and the reduction half-cell is termed as cathode.

At anode, oxidation occurs

2Na(s)2Na+(aq)+2e{\rm{2Na}}\left( s \right) \to 2{\rm{N}}{{\rm{a}}^ + }\left( {aq} \right) + 2{e^ - }\;\;

At cathode, reduction occurs,

Cl2(g)+2e2Cl(aq){\rm{C}}{{\rm{l}}_2}\left( g \right) + 2{e^ - } \to 2{\rm{C}}{{\rm{l}}^ - }\left( {aq} \right)

Ans:

The balanced half-cell reactions are:

Oxidation half reaction: 2Na(s)2Na+(aq)+2e{\rm{2Na}}\left( s \right) \to 2{\rm{N}}{{\rm{a}}^ + }\left( {aq} \right) + 2{e^ - }\;\;

Reduction half reaction: Cl2(g)+2e2Cl(aq){\rm{C}}{{\rm{l}}_2}\left( g \right) + 2{e^ - } \to 2{\rm{C}}{{\rm{l}}^ - }\left( {aq} \right)